whats the equation for the hyperbola described. Foci (3,7) and (7,7); vertex at (6,7)
Answer by lwsshak3(11628)

whats the equation for the hyperbola described. Foci (3,7) and (7,7); vertex at (6,7)
**
Given data shows this is a hyperbola with horizontal transverse axis. Its standard form of equation:
(x-h)^2/a^2-(y-k)^2/b^2=1, (h,k)=(x,y) coordinates of center.
..
x-coordinate of center=(7+3)/2=5 (midpoint formula)
y-coordinate=7
center: (5,7)
..
length of transverse axis=2=2a (2*distance from vertex to center)
a=1
a^2=1
..
Foci
2c=4 (distance between foci)
c=2
c^2=4
..
c^2=a^2+b^2
b^2=c^2-a^2=4-1=3
...
Equation:
(x-5)^2/1-(y-7)^2/3=1